OpenStudy (anonymous):

Constant Velocity Model and Constant Acceleration Model Combined! A riffle is aimed at the centre of a large target 200m away. The initial speed of the bullet is 500m/s. a) How far below the centre does the bullet strike the target? b) To hit the centre of the target, the barrel must be at an angle above the line of sight to the centre of the target. Find the angle of elevation of the barrel.

OpenStudy (anonymous):

I did a) and found that the bullet is 78cm away from the centre, but I have no idea how to work out part b)

OpenStudy (anonymous):

a) t=200/500=0.4 and y=-1/2gt^2=-0.78m=-78cm a is right

OpenStudy (anonymous):